### v2

@@ Appendix B Broader Impacts
Appendix C Detailed Explanations of LLMs Architecture
We focus on the autoregressive Transformer-based LLM architecture (Vaswani et al.,, 2017) based on prior works (Geva et al.,, 2021; Elhage et al.,, 2021; Geva et al.,, 2022; Dai et al.,, 2022; Meng et al., 2022a, ; Meng et al., 2022b, ; Yuksekgonul et al.,, 2024) with simplifications in certain explanations. Given an input prompt containing $T$ tokens $(t_{1},\ldots,t_{T})$ and each token $t_{i}$ belonging to a vocabulary $\mathcal{V}$, tokens are initially encoded by $d$-dimensional vectors $\mathbf{x}_{i}^{0}\in\mathbb{R}^{d}$ using an embedding matrix $W_{E}\in\mathbb{R}^{|\mathcal{V}|\times d}$.
The architecture has $L$ layers, and each layer consists of attention and MLP modules, which transform token embeddings to residual streams $(\mathbf{x}_{1}^{\ell},\ldots,\mathbf{x}_{T}^{\ell})\in\mathbf{X}^{\ell}$ at layer $\ell$, where $\mathbf{x}_{i}^{\ell}\in\mathbb{R}^{d}$. The residual stream at layer $\ell$ is a place where all attention and MLP modules at layer $\ell$ read from and write to (Elhage et al.,, 2021), and it is updated by the following equation for token $i$ at layer $\ell$:
@@ Appendix C Detailed Explanations of LLMs Architecture
The attention module mainly updates each token residual stream $\mathbf{x}_{i}^{l-1}$ by attending to all previous tokens in parallel. Specifically, the attention module contains $QK$ and $OV$ circuits, where the former operates $W_{Q},W_{K}\in\mathbb{R}^{d\times d}$ matrices and the latter operates $W_{O},W_{V}\in\mathbb{R}^{d\times d}$ matrices, respectively. Normally, $QK$ circuit determines the attention pattern $A^{\ell}$, i.e., where information is moved to and from the residual stream. $OV$ circuit further determines the attention output $\mathbf{a}_{i}^{\ell}$ based on the fixed attention pattern, i.e., what information is from the previous tokens’ position to the current token position (Elhage et al.,, 2021):
@@
where $\gamma(\cdot)$ is activation function, $W_{\text{in}}^{\ell}\in\mathbb{R}^{d\times d_{\text{m}}}$, and $W_{\text{out}}^{\ell}\in\mathbb{R}^{d_{\text{m}}\times d}$. $d_{\text{m}}$ is the dimensionality of MLP module, which is larger than $d$. MLP module is normally treated as key-value memories (Geva et al.,, 2021; Elhage et al.,, 2021; Geva et al.,, 2022; Dai et al.,, 2022), where columns of $W_{\text{in}[:,i]}^{\ell}$ and rows of $W_{\text{out}[i,:]}^{\ell}$ are viewed as keys and values, respectively. Given the input $\mathbf{x}_{i}^{\ell-1}$, the keys of MLP produce a vector of cofficients $\mathbf{k}_{i}^{\ell}=\gamma(W_{\text{in}}^{\ell}\mathbf{x}_{i}^{\ell-1})\in\mathbb{R}^{d_{\text{m}}}$, and they weights the corresponding values $\mathbf{v}_{i}^{\ell}$ in $W_{\text{out}}^{\ell}$. Therefore, Eq. 14 can be reformatted as [^7] We omit all bias $b_{Q},b_{K},b_{O},b_{V},b_{\text{in}},b_{\text{out}},b_{\text{U}}$ for simplification.:
@@
Appendix D Details of Each Dataset and Experimental Settings
  • Indirect Object Identification (IOI) Wang et al., (2022): IOI is a manually synthesised corpus used to understand a specific natural language task, where sentences such as “Afterwards Lisa and Rachel went to the garden, and Lisa gave a bone to” should be followed with “Rachel”. When two names are included in such sentences, the predicted name should not be the subject of the last clause. This task has been verified that GPT2 family works well (Wang et al.,, 2022; Merullo et al.,, 2024). However, we found that GPT2 family immediately fails this task if the input prompt is formatted as MCQ, where the incorrect subject of the last clause is placed in the ‘A’ choice.
  • Greater-than task (Greater) Hanna et al., (2024): Greater-than task is also a manually synthesised corpus, which is used to evaluate GPT2’s ability to sentences such as “The war lasted from the year 1732 to the year 17”, and model will predict valid two-digit end years, i.e., years > 32. However, we found the same anchored bias like IOI when this task is formatted as MCQ in Fig. 1, and GPT2 family also fails to predict valid years and the prediction is always anchored at incorrect choice ‘A’.
  • Logical Deduction of the Big-Bench (LD) [^8] https://github.com/google/BIG-bench/blob/main/bigbench/benchmark_tasks/logical_deduction/three_objects/task.json Srivastava et al., (2023): LD is a subtask which evaluates three-object logical deduction tasks, and it is used to measure whether model can parse information about multiple choices and their mutual relationships. Each MCQ in the task includes three similar objects in a naturally ordered context (e.g., books of various colours sitting on a shelf) and a set of simple clues regarding their placement (e.g., "the red book is to the right of the green book") such that no clue is redundant. The challenge is to assign the highest probability to correct MCQ choice about which object lies at which position.
  • ARC-Challenge (ARC) Clark et al., (2018): ARC is a real grade-school level, multiple-choice science questions, which includes Easy and Challenge versions. We choose the Challenge version to evaluate the anchored bias.
  • CommensenseQA (CSQA) Talmor et al., (2019): CSQA is a new multiple-choice question answering dataset that requires different types of commonsense knowledge to predict the correct answers.
@@ Appendix D Details of Each Dataset and Experimental Settings
Appendix E The Statistic Information of Each Test Dataset for GPT2 Models
@@
Appendix F Top-10 Unembeded Tokens from Identified Value Vectors
@@
Appendix G Top-10 Unembedded Tokens from Updated Value Vectors
@@
Appendix H Ablation Study of Different $\lambda_{2}$ in Eq. 8
We conduct an ablation study by changing $\lambda_{2}$ from 8 to 2 in Eq. 8, and evaluate the classification accuracy of each Infer. and Eva. dataset using GPT2 family. Table 10,11,12,13,14,15,16 show the ablation study results under different $\lambda_{2}$. We can find that the classification accuracy shows a decreasing trend with $\lambda_{2}$ from 8 to 2, which indicates that the effect of direct value vector update using Eq. 8 decreases.
@@
Appendix I Top-1 Classification Accuracy Changes Using Attention Pattern Recalibration
@@
Appendix J One-shot and Two-shot MCQ Classification Results
@@
Appendix K Damage to Updated GPT2 Family’s Performance on IOI and Greater Datasets
Based on the discussion in § 6, we further evaluate the damage of direct value vector update from MLP for the general ability of GPT2 family. We choose original IOI and Greater-than datasets as they have been verified that GPT2 family works well (Wang et al.,, 2022; Merullo et al.,, 2024; Hanna et al.,, 2024). Fig. 9 and Fig. 10 show that classification accuracy is acceptable when a value vector with a specific layer and dimensionality is updated, which matches the findings from Gu et al., (2024).
@@
Appendix L MCQ Prompt Template with Random Words
@@
Appendix M Full Circuit of Anchored Bias of Each GPT2 Model

### v3

@@ Appendix A Broader Impacts
Appendix B Detailed Explanations of LLMs Architecture
We focus on the autoregressive Transformer-based LLM architecture (Vaswani et al., 2017) based on prior works (Geva et al., 2021, 2022; Elhage et al., 2021; Dai et al., 2022; Meng et al., 2022, 2023; Yuksekgonul et al., 2024) with simplifications in certain explanations. Given an input prompt containing $T$ tokens $(t_{1},\ldots,t_{T})$ and each token $t_{i}$ belonging to a vocabulary $\mathcal{V}$, tokens are initially encoded by $d$-dimensional vectors $\mathbf{x}_{i}^{0}\in\mathbb{R}^{d}$ using an embedding matrix $W_{E}\in\mathbb{R}^{|\mathcal{V}|\times d}$.
The architecture has $L$ layers, and each layer consists of attention and MLP modules, which transform token embeddings to residual streams $(\mathbf{x}_{1}^{\ell},\ldots,\mathbf{x}_{T}^{\ell})\in\mathbf{X}^{\ell}$ at layer $\ell$, where $\mathbf{x}_{i}^{\ell}\in\mathbb{R}^{d}$. The residual stream at layer $\ell$ is a place where all attention and MLP modules at layer $\ell$ read from and write to (Elhage et al., 2021), and it is updated by the following equation for token $i$ at layer $\ell$:
@@ Appendix B Detailed Explanations of LLMs Architecture
The attention module mainly updates each token residual stream $\mathbf{x}_{i}^{l-1}$ by attending to all previous tokens in parallel. Specifically, the attention module contains $QK$ and $OV$ circuits, where the former operates $W_{Q},W_{K}\in\mathbb{R}^{d\times d}$ matrices and the latter operates $W_{O},W_{V}\in\mathbb{R}^{d\times d}$ matrices, respectively. Normally, $QK$ circuit determines the attention pattern $A^{\ell}$, i.e., where information is moved to and from the residual stream. $OV$ circuit further determines the attention output $\mathbf{a}_{i}^{\ell}$ based on the fixed attention pattern, i.e., what information is from the previous tokens’ position to the current token position (Elhage et al., 2021):
@@
where $\gamma(\cdot)$ is activation function, $W_{\text{in}}^{\ell}\in\mathbb{R}^{d\times d_{\text{m}}}$, and $W_{\text{out}}^{\ell}\in\mathbb{R}^{d_{\text{m}}\times d}$. $d_{\text{m}}$ is the dimensionality of MLP module, which is larger than $d$. MLP module is normally treated as key-value memories (Geva et al., 2021, 2022; Elhage et al., 2021; Dai et al., 2022), where columns of $W_{\text{in}[:,i]}^{\ell}$ and rows of $W_{\text{out}[i,:]}^{\ell}$ are viewed as keys and values, respectively. Given the input $\mathbf{x}_{i}^{\ell-1}$, the keys of MLP produce a vector of cofficients $\mathbf{k}_{i}^{\ell}=\gamma(W_{\text{in}}^{\ell}\mathbf{x}_{i}^{\ell-1})\in\mathbb{R}^{d_{\text{m}}}$, and they weights the corresponding values $\mathbf{v}_{i}^{\ell}$ in $W_{\text{out}}^{\ell}$. Therefore, Eq. 14 can be reformatted as [^8] We omit all bias $b_{Q},b_{K},b_{O},b_{V},b_{\text{in}},b_{\text{out}},b_{\text{U}}$ for simplification.:
@@
Appendix C Details of Each Dataset and Experimental Settings
  • Indirect Object Identification (IOI) Wang et al. (2023): IOI is a manually synthesised corpus used to understand a specific natural language task, where sentences such as “Afterwards Lisa and Rachel went to the garden, and Lisa gave a bone to” should be followed with “Rachel”. When two names are included in such sentences, the predicted name should not be the subject of the last clause. This task has been verified that GPT2 family works well (Wang et al., 2023; Merullo et al., 2024). However, we found that GPT2 family immediately fails this task if the input prompt is formatted as MCQ, where the incorrect subject of the last clause is placed in the ‘A’ choice.
  • Greater-than task (Greater) Hanna et al. (2023): Greater-than task is also a manually synthesised corpus, which is used to evaluate GPT2’s ability to sentences such as “The war lasted from the year 1732 to the year 17”, and model will predict valid two-digit end years, i.e., years > 32. However, we found the same anchored bias like IOI when this task is formatted as MCQ in Fig. 1, and GPT2 family also fails to predict valid years and the prediction is always anchored at incorrect choice ‘A’.
  • Logical Deduction of the Big-Bench (LD) [^9] https://github.com/google/BIG-bench/blob/main/bigbench/benchmark_tasks/logical_deduction/three_objects/task.json Srivastava et al. (2023): LD is a subtask which evaluates three-object logical deduction tasks, and it is used to measure whether model can parse information about multiple choices and their mutual relationships. Each MCQ in the task includes three similar objects in a naturally ordered context (e.g., books of various colours sitting on a shelf) and a set of simple clues regarding their placement (e.g., "the red book is to the right of the green book") such that no clue is redundant. The challenge is to assign the highest probability to correct MCQ choice about which object lies at which position.
  • ARC-Challenge (ARC) Clark et al. (2018): ARC is a real grade-school level, multiple-choice science questions, which includes Easy and Challenge versions. We choose the Challenge version to evaluate the anchored bias.
  • CommensenseQA (CSQA) Talmor et al. (2019): CSQA is a new multiple-choice question answering dataset that requires different types of commonsense knowledge to predict the correct answers.
@@ Appendix C Details of Each Dataset and Experimental Settings
Appendix D The Statistic Information of Each Test Dataset for GPT2 Models
@@
Appendix E Top-10 Unembeded Tokens from Identified Value Vectors
@@
Appendix F Top-10 Unembedded Tokens from Updated Value Vectors
@@
Appendix G Ablation Study of Different $\lambda_{2}$ in Eq. 8
We conduct an ablation study by changing $\lambda_{2}$ from 8 to 2 in Eq. 8, and evaluate the classification accuracy of each Infer. and Eva. dataset using GPT2 family. Table 11,12,13,14,15,16,17 show the ablation study results under different $\lambda_{2}$. We can find that the classification accuracy shows a decreasing trend with $\lambda_{2}$ from 8 to 2, which indicates that the effect of direct value vector update using Eq. 8 decreases.
@@
Appendix H Top-1 Classification Accuracy Changes Using Attention Pattern Recalibration
@@
Appendix I One-shot and Two-shot MCQ Classification Results
@@
Appendix J Damage to Updated GPT2 Family’s Performance on IOI and Greater Datasets
Based on the discussion in § 7, we further evaluate the damage of direct value vector update from MLP for the general ability of GPT2 family. We choose original IOI and Greater-than datasets as they have been verified that GPT2 family works well (Wang et al., 2023; Merullo et al., 2024; Hanna et al., 2023). Fig. 9 and Fig. 10 show that classification accuracy is acceptable when a value vector with a specific layer and dimensionality is updated, which matches the findings from Gu et al. (2024).
@@
Appendix K MCQ Prompt Template with Random Words
@@
Appendix L Full Circuit of Anchored Bias of Each GPT2 Model